OpenStudy (anonymous):

What is the sum of a 17-term arithmetic sequence where the first term is 7 and the last term is -105?

OpenStudy (anonymous):

whoevers first

OpenStudy (cwrw238):

the formula for this is Sum = ( n/2){ a + l) whre n = number of terms , a = first term and l = last term

OpenStudy (anonymous):

(-105-7)/16=-7(common difference) (-105+7)(-105-7+7)/(-14)=-810

OpenStudy (anonymous):

-833 -893 -913 -943 these are the choices

OpenStudy (cwrw238):

= 8.5 *(7 - 105)

OpenStudy (cwrw238):

= -833

OpenStudy (anonymous):

833, but i got a positvie

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yes i had a mistake. sorry :)